How much of a 80% alcohol solution must be mixed with 60 gallons of a 24% alcohol solution to obtain a solution that is 70% alcohol?

Answer by checkley71(8403)

.8x+.24*60=.7(x+60)
.8x+14.4=.7x+42
.8x-.7x=42-14.4
.1x=27.6
x=27.6/.1
x=276 gallons of 80% alcohol is required.
proof
.8*276+14.4=.7(276+60)
220.8+14.4=.7*336
235.2=235.2